# Spatiotemporal Change-Points in Development Discourse: Insights from Social Media in Low-Resource Contexts

### v1

@@ 5.4. Crisis Communication and Informative Content
The challenge of identifying informative content amid high-volume social media streams during crises has been well-documented (Ghafarian and Yazdi, 2020). As Ghafarian and Yazdi note, crisis communications exhibit increasing complexity that complicates automated classification. Our qualitative coding approach addresses this by grounding topic interpretation in domain expertise, ensuring that the development-related discourse we identify reflects genuine public concerns rather than noise. Earlier work has shown that the ”long now” of crises is not well understood (Soden and Palen, 2018; Dailey, 2020), and that online spaces complement discourse in other areas of people’s lives as they deal with long-term effects of disaster (Randazzo et al., 2025). Our findings extend this insight by showing that development interventions such as infrastructure projects can generate lasting shifts in public discourse, extending the ”long now” phenomenon beyond disaster contexts.
@@ 6. Limitations and Future Work
Figure 4. The Dual-Mode WhatsApp Sensing Framework. This hierarchical tree structure facilitates efficient information dissemination from policymakers to local communities (downward flow) while enabling bidirectional engagement through passive sensing (topic modeling of donated messages) and active sensing (deployment of polls and surveys)
@@ 7. Conclusion
This study demonstrates that geotagged social media reveal the evolving contours of development discourse in low-resource contexts. Using BERTopic, PELT, and qualitative coding on Zambian X data, we identified seven key themes and critical change-points linked to events such as COVID-19 and a geothermal project. We introduced durable discourse to capture the persistence of development concerns over time. Our findings show that social media can act as a real-time sensor of policy flashpoints, supporting more responsive and equitable development planning. They also underscore the need for cross-platform, multimodal, and participatory approaches to address data gaps and platform biases.

### v2

@@ 5.6. Crisis Communication and Informative Content
The challenge of identifying informative content amid high-volume social media streams during crises has been well-documented (Ghafarian and Yazdi, 2020). As Ghafarian and Yazdi note, crisis communications exhibit increasing complexity that complicates automated classification. Crisis informatics research has established that disasters produce temporal phases with distinct communication needs: the acute phase demands rapid information dissemination, while recovery phases require sustained community engagement (Reuter and Kaufhold, 2018; Palen and Anderson, 2016). Our qualitative coding approach addresses this by grounding topic interpretation in domain expertise, ensuring that the development-related discourse we identify reflects genuine public concerns rather than noise.
Stieglitz et al. (Stieglitz et al., 2018) demonstrate how social media enables collective sense-making during extreme events, with different actors contributing to shared understanding through distinct roles. Extending this, Mirbabaie et al. (Mirbabaie et al., 2020) introduce the complementary concept of “sense-giving”—the intentional provision of information by authoritative actors to shape collective meaning creation. Their analysis of Hurricane Harvey communications reveals how organizations can maintain influence over crisis discourse through sustained, strategic communication. This sense-making/sense-giving dynamic has direct implications for understanding durable discourse: persistent engagement may reflect ongoing sense-giving efforts by authoritative actors, ongoing collective sense-making by affected communities, or both.
Earlier work has shown that the “long now” of crises is not well understood (Soden and Palen, 2018; Dailey, 2020), and that online spaces complement discourse in other areas of people’s lives as they deal with long-term effects of disaster (Randazzo et al., 2025). Randazzo et al.’s (Randazzo et al., 2025) analysis of Hurricane Ida recovery discourse over eight months demonstrates that crises have both acute and chronic effects on communities, with online and offline discourse showing significant thematic overlap around long-term consequences. Similarly, research on disaster response trajectories (De Silva et al., 2023) identifies prototypical stages through which community responses evolve on social media, including a “honeymoon phase” of social cohesion followed by a “disillusionment phase” as communities confront long-term repercussions.
Our findings extend this insight by showing that development interventions such as infrastructure projects can generate lasting shifts in public discourse, extending the “long now” phenomenon beyond disaster contexts to planned development. The geothermal project in Southern Province exemplifies how infrastructure development can trigger durable discourse patterns that persist through multiple recovery-like phases, requiring sustained rather than episodic policy engagement.
@@ 5.7. Toward Participatory Development Sensing
Our findings also have implications for the design of civic technologies that aim to surface and sustain citizen voice. DiSalvo et al. (DiSalvo et al., 2014) argue that HCI design can express “matters of concern”—issues that gather publics around shared problems. Our durable discourse framework provides a measurement approach for identifying when citizen concerns achieve this gathering function versus remaining ephemeral. Development practitioners designing participatory sensing systems could use change-point analysis to distinguish between short-term attention spikes (requiring immediate response) and durable discourse patterns (warranting sustained policy engagement). This aligns with DIS scholarship on community-driven design (Hekler et al., 2019), which emphasizes examining assumptions behind design methods to increase likelihood of positive social impact.
Furthermore, Le Dantec and DiSalvo (Dantec and DiSalvo, 2013) describe how participatory design “infrastructures” the formation of publics—creating the conditions for collective attention to emerge around shared concerns. Our spatiotemporal methodology complements this perspective by measuring when such publics coalesce and how long they persist. The distinction between ephemeral COVID-19 discourse and durable infrastructure discourse in our findings illustrates different modes of public formation: crisis-driven versus sustained structural engagement. Recent DIS work on citizen sensing (Prandi et al., 2025) similarly emphasizes how interactive systems can be designed to transform data collection into meaningful civic participation—a goal our temporal measurement framework directly supports by identifying which citizen-generated signals warrant sustained organizational attention.
@@ 6. Limitations and Future Work
Figure 4. The Dual-Mode WhatsApp Sensing Framework. This hierarchical tree structure facilitates efficient information dissemination from policymakers to local communities (downward flow) while enabling bidirectional engagement through passive sensing (topic modeling of donated messages) and active sensing (deployment of polls and surveys).
@@ 7. Conclusion
This study demonstrates that geotagged social media reveal the evolving contours of citizen voice in development contexts. Using BERTopic, PELT, and qualitative coding on Zambian X data, we operationalized durable discourse—sustained public engagement that persists beyond event-driven attention spikes—and distinguished it from ephemeral crisis-driven discourse. Our findings show that social media platforms function as networked public squares where development concerns are articulated, contested, and sustained over time. By measuring the temporal persistence of citizen voice through explicit criteria (temporal persistence, thematic coherence, and structural signature), our framework contributes to digital civics scholarship and informs the design of participatory development monitoring systems. Future work should extend this approach across platforms, particularly privacy-preserving analysis of WhatsApp and other encrypted messaging services that dominate communication in the Global South.